Would someone please show me step by step how to figure out this problem?
Find the slope
-6x-9y=12

Found 2 solutions by stanbon, scott8148:
Answer by stanbon(75887)

Find the slope
-6x-9y=12
-----------
You need to put the equation into slope-intercept form: y=mx+b
Solve for y:
9y=-6x-12
y = (-6/9)x-(12/9)
y = (-2/3)x - (4/3)
The slope is -2/3
============
Cheers,
Stan H.

Answer by scott8148(6628)

the general form of the slope-intercept equation is y=mx+b, where m is the slope and b is the y-intercept

you have to use algebra to get y by itself on one side of the equation, then the coefficient of the x term is the slope

-6x-9y=12 ... adding 6x to both sides gives -9y=6x+12 ... dividing the equation by -9 gives y=-(2/3)x-(4/3)

so the slope is -2/3 ... and the y-intercept is -4/3